# Towards Hierarchical Multi-Agent Workflows for Zero-Shot Prompt Optimization

### v1

@@ 5.1 Main Evaluation
All experiments in the main evaluation are conducted using the open-source model ‘Mixtral-8x7B-v0.1’ [8] as the LLM agent.
@@ 5.2 Further Analysis
A case study. We conduct qualitative analysis of the the prompts generated in our workflow and how it helps the final response generation procedure. In Fig. 5, we illustrate the hierarchical process of generating responses using HMAW. The process begins with the CEO layer, which sets high-level guidelines tailored to the user’s background and specific inquiry. The Manager layer then interprets these guidelines, adding detail and context to direct the response formulation. Finally, the Worker layer produces the final response, ensuring it is coherent, professional, and empathetic, taking into full account the user’s limited familiarity with the topic. This structured approach ensures that the response effectively addresses the user’s query and matches their level of understanding. More case studies for other datasets are available in the Appendix Section C.
@@ Appendix C More Case Studies
Limitation
Our method is limited in deployment environments where computational resources are extremely constrained because it requires more than one time of inference for the LLM.
Impact statement
This paper presents work aimed at advancing the field of multi-agent cooperation. There are many potential societal consequences of our work, such as improved human-computer interaction.

### v2

@@ 5.1 Main Evaluation
All experiments in the main evaluation are conducted using the open-source model ‘Mixtral-8x7B-v0.1’ (Jiang et al., 2024) as the LLM agent.
@@ 5.2 Further Analysis
A case study. We conduct qualitative analysis of the the prompts generated in our workflow and how it helps the final response generation procedure. In Fig. 5, we illustrate the hierarchical process of generating responses using HMAW. The process begins with the CEO layer, which sets high-level guidelines tailored to the user’s background and specific inquiry. The Manager layer then interprets these guidelines, adding detail and context to direct the response formulation. Finally, the Worker layer produces the final response, ensuring it is coherent, professional, and empathetic, taking into full account the user’s limited familiarity with the topic. This structured approach ensures that the response effectively addresses the user’s query and matches their level of understanding.